The following information summarizes the standard cost for producing one metal tennis racket frame at Spaulding Industries. In addition, the variances for one month's production are given. Assume that all inventory accounts have zero balances at the beginning of the month.

Standard Cost Per Unit Standard Monthly Costs
Materials $ 4.00 $ 8,400
Direct Labor 2 hrs. @ $2.60 5.20 10,920
Factory Overhead:
Variable 1.80 3,780
Fixed 5.00 10,500
$ 16.00 $ 33,600
Variances:
Material price 244.75 unfavorable
Material quantity 500.00 unfavorable
Labor rate 520.00 favorable
Labor efficiency 2,080.00 unfavorable

What were the actual direct labor hours worked during the month?
Answer #1

Number of units = $33,600/16.00 = 2,100

Labor efficiency = $2,080 U = [AH - (2,100 x 2)] x $2.60

$2,080 = [AH - (2,100 x 2)] x $2.60

$2,080 / $2.60 = AH - (2,100 x 2)

800 = AH - 4,200

AH = 800 + 4,200

AH = 5,000